When 4 is added to a certain number the result is the same as halving the number and adding 10
Answer by fractalier(6550)

Call the number x. We can then write
x + 4 = (1/2)x + 10
Now subtract half x and 4 from both sides and we get
(1/2)x + 4 = 10
(1/2)x = 6
Now multiply by two and get
x = 12